# Hypercube-RAG: Hypercube-Based Retrieval-Augmented Generation for In-domain Scientific Question-Answering

### v1

@@ 4.2 Baselines
We select three types of methods for comparison: 1) sparse retriever BM25 Robertson and Walker (1994); 2) dense embedding retrieval methods, Contriever Izacard et al. (2022), e5 Wang et al. (2024a), Nvidia/NV-Embedv2 Lee et al. (2025); 3) graph-based methods, GraphRAG Edge et al. (2024), LightRAG Guo et al. (2024), and HippoRAG Gutiérrez et al. (2024) and HippoRAG 2 Gutiérrez et al. (2025). All experiments were conducted on an A100 GPU with 80 GB memory. More details are presented in Appendix D.
| Method | Hurricane | Geography | Aging Dam | |||||||||
| Automatic Metric | LLM-as-Judge | Automatic Metric | LLM-as-Judge | Automatic Metric | LLM-as-Judge | |||||||
| BLEU | BERTScore | Correct | Complete | BLEU | BERTScore | Correct | Complete | BLEU | BERTScore | Correct | Complete | |
| No Retrieval | ||||||||||||
| GPT-4o | 2.6 | 66.8 | 57.1 | 23.5 | 2.8 | 63.9 | 64.2 | 42.1 | 5.5 | 62.5 | 56.9 | 19.6 |
| DeepSeek-R1 | 2.2 | 69.2 | 53.4 | 28.5 | 1.3 | 61.7 | 57.9 | 45.7 | 8.7 | 71.6 | 35.4 | 18.8 |
| Qwen2.5-7B-Instruct | 1.2 | 46.9 | 23.5 | 6.9 | 1.7 | 47.1 | 32.1 | 12.6 | 1.9 | 43.4 | 15.5 | 6.5 |
| Llama-3.3-70B-Instruct | 2.0 | 60.2 | 39.8 | 14.6 | 1.6 | 51.3 | 49.3 | 21.2 | 2.4 | 46.7 | 39.7 | 17.1 |
| Llama-4-Scout-17B-16E-Instruct | 2.3 | 62.7 | 53.9 | 39.1 | 3.1 | 61.3 | 48.8 | 21.9 | 5.2 | 64.2 | 34.4 | 15.1 |
| Sparse and Dense Embedding Retrieval | ||||||||||||
| BM25 Robertson and Walker (1994) | 10.5 | 77.5 | 74.3 | 60.8 | 7.9 | 73.3 | 79.7 | 70.8 | 18.1 | 80.2 | 85.7 | 75.9 |
| Contriever Izacard et al. (2021) | 11.3 | 78.6 | 75.7 | 60.6 | 7.5 | 74.1 | 82.5 | 73.6 | 16.7 | 76.8 | 70.9 | 60.9 |
| e5 Wang et al. (2024a) | 11.9 | 79.1 | 78.3 | 63.9 | 7.1 | 73.9 | 81.2 | 71.2 | 19.7 | 79.8 | 88.9 | 75.5 |
| NV-Embed Lee et al. (2025) | 12.2 | 79.1 | 76.9 | 59.6 | 7.5 | 75.4 | 81.9 | 72.3 | 18.8 | 79.6 | 79.4 | 73.6 |
| Graph-based RAG | ||||||||||||
| LightRAG Guo et al. (2024) | 7.6 | 72.2 | 64.1 | 47.6 | 7.0 | 67.8 | 69.8 | 59.6 | 12.8 | 75.4 | 65.9 | 46.5 |
| GraphRAG Edge et al. (2024) | 6.5 | 72.8 | 68.8 | 50.9 | 6.9 | 69.4 | 70.6 | 63.8 | 11.8 | 72.4 | 72.5 | 58.6 |
| HippoRAG Gutiérrez et al. (2024) | 6.8 | 73.6 | 70.5 | 48.8 | 7.8 | 70.3 | 76.9 | 69.7 | 12.4 | 73.8 | 80.9 | 58.6 |
| HippoRAG 2 Gutiérrez et al. (2025) | 8.6 | 76.3 | 86.7 | 53.5 | 9.2 | 74.6 | 80.6 | 67.5 | 11.5 | 74.1 | 88.2 | 53.7 |
| Our Method | ||||||||||||
| Hypercube-RAG | 13.2 | 82.5 | 80.6 | 66.2 | 9.3 | 76.8 | 83.6 | 74.8 | 21.5 | 82.6 | 86.9 | 77.8 |
4.3 Evaluation Metrics
@@
Following the work Cui et al. (2025), we evaluate the quality of LLMs’ responses using automated metrics derived from token-level representations, including BLEU Papineni et al. (2002) and BertScore Zhang* et al. (2020) to provide the standard assessment of response quality. Since the settings are open-text responses, we also employ LLM-as-a-Judge that assesses correctness and completeness. All LLM-based evaluations use GPT-4o as the judge. The prompts used are in Appendix A.
5 Results and Analysis
5.1 Accuracy
@@
We compare our Hypercube-RAG with semantic RAG methods, graph-based RAG methods, and LLMs without retrieval. Table 2 reveals three key findings. First, Hypercube-RAG consistently outperforms other baselines across three datasets, demonstrating its effectiveness in enhancing the capabilities of LLMs for in-domain scientific question-answering. Compared to the second-best method, Hypercube-RAG improves the accuracy by measured as relative gains across three of four metrics used. Second, all RAG methods perform much better than direct inference of LLMs without retrieval, as expected, showing the effectiveness of incorporating external knowledge for in-domain QA tasks. Finally, we also evaluate Hypercube-RAG against a range of baseline methods across multiple LLMs developed by different companies. As illustrated in Figure 5, our method achieves superior performance across all models regardless of the underlying LLM architectures.
5.2 Efficiency
@@
We study the response speed of diverse RAG methods over the corpus size and report the retrieval time for one query in Table 3. We break down our analysis as follows. Within corpus size . BM25 achieves the fastest retrieval among all evaluated methods without surprise. In contrast, semantic and graph-based RAGs incur significantly higher retrieval costs. This is particularly pronounced for graph-based methods due to the computational burden of search paths in large-scale graphs, posing scalability challenges. Notably, our Hypercube-RAG substantially reduces the retrieval time by one to two orders of magnitude compared to both semantic and graph-based methods, underscoring the effectiveness of the hypercube structure in optimizing retrieval efficiency. We attribute this to the retrieval with compact cube labels introduced in Section 3.3. Although it is slightly slower than BM25, the marginal increase in retrieval time is a worthwhile trade-off for significantly improved accuracy, as shown in Table 2. Beyond corpus size . To mimic the real-world scenario that the external knowledge database is usually large and noisy, we expand the hurricane corpus from 844 to 11,539 documents (around 14 times). The added documents are related to other topics, such as “Pollution” or “Ozone Holes”, introducing additional noise. Hypercube-RAG maintains the lowest retrieval time per query (see the last column in Table 3), highlighting the efficiency of the cube label-based retrieval mechanism. This is likely because most noisy documents result in empty or sparsely populated cube cells, as they lack corresponding in-domain fine-grained labels pre-defined in our hypercube. As a result, despite a significantly expanded corpus, the hypercube can efficiently bypass these irrelevant documents. In contrast, other baseline methods exhibit greater sensitivity to corpus size, resulting in increased retrieval times.
@@
In addition, Table 4 shows that the accuracy of the evaluated methods drops due to the inclusion of these noisy off-topic documents. However, Hypercube-RAG still outperforms others, demonstrating robust resilience to noisy data.
| Methods | BERTScore () | ||
@@
5.3 Explainability
The retrieval process with our hypercube is inherently explainable through its associated cube labels. Considering the query “How much rainfall did Melbourne Beach, Florida receive from Tropical Storm Fay?”, illustrated in Figure 4, Hypercube-RAG returns relevant documents by matching the decomposed components with fine-grained labels, ‘Melbourne Beach’, ‘Tropical Storm Fay’, and ‘rainfall’, along hypercube dimensions: LOCATION, EVENT, and THEME. Table 5 shows that three retrieved documents (Doc. 565, 246, 535), which are represented with fine-grained cube labels. These cube labels clearly explain why those documents are returned. For example, Doc. 565 is ranked the first to return, since it contains all three key components of the query. The detailed document contents are included in Appendix G.
@@ 5.3 Explainability
5.4 Ablation Study
@@
As shown in Section 3.3, Hypercube-RAG combines two retrieval strategies: sparse exact matching and dense embedding retrieval. Section 3.4 introduces the document ranking algorithm used to filter and prioritize the retrieved results. To study their effectiveness, we conduct an ablation study by removing each of them from Hypercube-RAG, which are represented as No-Sparse, No-Dense, and No-Ranking, respectively. Figure 7 shows that the full version of Hypercube-RAG consistently outperforms all other variants, verifying positive contributions of each constituent. More specifically, it justifies the benefit of coupling these two strategies for in-domain theme-specific question-answering tasks, and the necessity of the ranking algorithm to filter possible “noisy” documents. The ablation study on hypercube dimensions is in Appendix F.
5.5 Parameter Study
Hypercube-RAG combines sparse exact match and semantic retrieval strategies, as introduced in Section 3.3. We investigate the effect of the semantic retrieval strategy on answer quality under varying similarity thresholds, denoted by . Figure 8 shows that performance improves with increasing threshold values when , suggesting that higher similarity scores are more effective for retrieving semantically relevant documents. However, when , the behavior of semantic RAG begins to gradually degrade to the sparse exact matching strategy, resulting in a decline in performance. This is primarily due to overly strict matching criteria, which reduce the number of retrieved documents and thus limit the available context.
5.6 Case Study
An additional case study in Figure 6 presents that both Semantic-RAG and HippoRAG-2 fail to generate the correct answer due to inaccurate document retrieval. These methods tend to prioritize returning documents that are semantically similar in the embedding space. While the retrieved documents mention general topics such as above-normal, near-normal, and below-normal hurricane season, they miss the specific location ‘Atlantic’ and organization ‘Climate Prediction Center’. In contrast, our Hypercube-RAG returns the correct answer. It accurately identifies and retrieves documents containing key elements of the query, including ‘Atlantic’, ‘hurricane season’, and ‘Climate Prediction Center’, corresponding to the hypercube dimensions, LOCATON, THEME, and ORGANIZATION, respectively. This retrieval approach intuitively prioritizes specific contextual information over abstract, ambiguous references such as above-normal, near-normal, and below-normal alone.
6 Conclusions
In this work, we start by analyzing the strengths and limitations of existing RAG methods for in-domain theme-specific applications. To retain their advantages while addressing key challenges, we propose Hypercube-RAG, innovating retrieval based on fine-grained labels with a multidimensional structure (i.e., hypercube). The label-based retrieval integrates sparse lexical and dense semantic strategies, making it accurate, efficient, and inherently explainable. Empirical results demonstrate that our method consistently outperforms existing RAG baselines in accuracy and efficiency, while providing interpretable provenance.
Limitations
The current hypercube construction partially relies on manual input from domain experts to predefine different dimensions. To enhance generalizability across diverse scientific question-answering tasks, we aim to dynamically design the hypercube dimensions. One possibility is to leverage large language models (LLMs) to automate this process.
Ethics Statement
We conduct our experiments on synthesized QA datasets collected from a publicly available scientific corpus. Our proposed approach and findings do not pose any foreseeable harm to individuals or groups. Overall, we do not anticipate any significant ethical concerns with this work.
Acknowledgments
This work is supported by the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE), which is funded by the National Science Foundation (NSF) under award number: 2118329. Any opinions, findings, conclusions, or recommendations expressed herein are those of the authors and do not necessarily represent the views of NSF.
Appendix
Appendix A Prompt Templates
Appendix B QA Categories
Appendix C Discussion
C.1 Why does Hypercube-RAG outperform other RAG methods?
Our Hypercube-RAG is particularly proposed for domain-focused analysis, where domain-specific information, e.g., location, date, event, event, theme, is central to the query. In such cases with theme-specific information, our hypercube enables precise and efficient retrieval by accessing only one or a few cells in a theme-specific hypercube, as each cell encapsulates multiple dimensions of information associated with a document. These fine-grained document labels can be represented in a hypercube along different dimensions. When a query comes in, accessing one or a few hypercubes presents a higher chance of retrieving highly relevant documents covering all key information contained in a query - Accuracy. Additionally, retrieval is operated with the cube labels significantly speeds up the retrieval time since those labels are in a compact space - Efficiency. In contrast, sparse exact match and dense embedding retrieval require traversing all documents and contents inside. Moreover, the retrieval process is inherently explainable, as the searching is based on the document’s labels represented by a hypercube - Explainability.
Case 1: simple query.
How much rainfall did Melbourne Beach, Florida receive from Tropical Storm Fay? The relevant documents can be retrieved from one cube cell in a hypercube <Melbourne Beach, rainfall, Tropical Storm Fay >.
Case 2: long query with multiple topics.
In cases where a query is very diverse, they may need to access multiple cube cells such that the query information can be covered as much as possible. For example, the diverse query could be “What consequences were caused by Tropical Storm Fay, such as how much rainfall Melbourne Beach and Palm Bay in Florida received, and how much monetary losses were caused?”. Those cube cells needed could be <Melbourne Beach, rainfall, Tropical Storm Fay >, <Palm Beach, rainfall, Tropical Storm Fay >, <Melbourne Beach, monetary losses, Tropical Storm Fay >, <Palm Beach, monetary losses, Tropical Storm Fay>.
Case 3: diverse query with shifted topics.
Given a more complicated query including shifted topics, e.g., “How did the monetary loss caused by Tropical Storm Fay impact the industrial layoff in Miami Downtown?”, cube cells across two hypercubes are needed <Miami Downtown, Industrial Layoff, Tropical Storm Fay > and <Miami Downtown, Monetary loss, Tropical Storm Fay > to get more relevant documents.
C.2 Why did we use in-domain scientific QA datasets rather than general ones?
According to the above discussion, hypercube construction is sensitive to themes in various domains. While similar or related themes are allowed to be included in a hypercube, each theme-specific dataset should have a separate hypercube to maximize its power. However, the general QA datasets, e.g., PopQA Mallen et al. (2022) and NQ Kwiatkowski et al. (2019), typically present articles across multiple themes. The process of constructing multiple hypercubes will require intensive efforts (e.g., design hypercube schema for different datasets.), which is out of the scope of this study. This is why we particularly focus on in-domain theme-specific QA datasets.
Appendix D Implementation Details
We selected three types of RAG baseline methods for comparison: sparse retriever BM25, dense embedding retrieval, and graph-based methods.
For sparse retriever BM25, we implement it using a GitHub repository444https://github.com/dorianbrown/rank_bm25.
For most experiments, we used the GPT-4o555https://platform.openai.com/docs/models/gpt-4o as the LLM base unless explicitly specified. In Table 2, we also include baseline methods without RAG, such as DeepSeek-R1666https://github.com/deepseek-ai/DeepSeek-R1, Qwen2.5-7B-Instruct777https://qwenlm.github.io/blog/qwen2.5/, Llama-3.3-70B-Instruct888https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct, and Llama-4-Scout-17B-16E-Instruct999https://huggingface.co/meta-llama/Llama-4-Scout-17B-16E-Instruct. Note that DeepSeek-R1 is a reasoning model, we set 0.6 and 1 as the temperature by following their recommendations. Since its output also includes the reasoning/thinking process, we exclude it and consider the final output as the answer. The temperatures for other models are set to 0.
In the evaluation phase, we compute BLEU scores and BERTScores with the NLTK101010https://www.nltk.org/ and SentenceTransformer111111https://sbert.net/ Python packages, respectively. We also use an LLM (GPT-4o) as the judge to compute the correctness and comprehensiveness (see the prompt in Appendix A).
Appendix E Details on Theme-specific Extraction
A small number of specialized terms are manually added, such as “hurricane landfall”, “hurricane hitting”, “hurricane intensity”, “sustained winds”.
Appendix F Ablation Study on Hypercube Dimensions
@@
We conduct an ablation study on the hypercube dimensions: Location, Event, Date, Organization, Person, and Theme. To study their effectiveness, we conduct an ablation study by removing each of the dimensions, which are represented as No-Location, No-Event, No-Date, No-Organization, No-Person, and No-Theme, respectively. Table 6 shows that the full version of Hypercube-RAG consistently outperforms all other variants, verifying positive contributions of each dimension. More specifically, we observe that the performance drops the most when removing the Location and Theme dimensions, justifying the high importance of these two hypercube dimensions for in-domain theme-specific question-answering.

### v2

@@ 4.2. Baselines
We select four types of methods for comparison: 1) sparse retriever BM25 (Robertson and Walker, 1994); 2) dense embedding retrieval methods: Contriever (Izacard et al., 2022), e5 (Wang et al., 2024b), Nvidia/NV-Embedv2 (Lee et al., 2025); 3) graph-based methods: GraphRAG (Edge et al., 2024), LightRAG (Guo et al., 2024), and HippoRAG (Gutiérrez et al., 2024) and HippoRAG 2 (Gutiérrez et al., 2025); and (4) LLMs without retrieval. We implement the sparse BM25 retriever using a GitHub repository444https://github.com/dorianbrown/rank_bm25. For dense embedding retrievers, we utilize their models uploaded to Hugging Face. Experiments for graph-based baselines using their official GitHub repositories. All experiments were conducted on a single A100 GPU with 80 GB memory.
4.3. Evaluation Metrics
@@
We utilize multiple metrics for a comprehensive evaluation of different methods. To measure the QA performance, following the work (Cui et al., 2025), we evaluate the quality of LLMs’ responses using automated metrics derived from token-level representations, including F1 and Semantic score (Chandrasekaran and Mago, 2021) to provide the standard assessment of response quality. Since the settings are open-text responses, we also employ LLM-as-a-Judge that assesses correctness and completeness. All LLM-based evaluations use GPT-4o as the judge. The prompts and other parameters used are in Appendix B. To measure the QA performance, following the HippoRAG work (Gutiérrez et al., 2025), precison@ and recall@ rates are computed to measure the retrieval performance.
5. Results
5.1. Accuracy
5.1.1. QA performance.
@@ 5.1. Accuracy
Table 4.1 reports the results for QA performance. The observations and findings are as follows. We noticed that the state-of-the-art baselines present varying performance over different datasets. Most of them perform better on SciFact than the other two datasets. Notably, Hypercube-RAG consistently outperforms other baselines across all three datasets, demonstrating its effectiveness in enhancing the capabilities of LLMs for scientific question-answering tasks. Hypercube-RAG surpasses the second-best models with an average improvement by on F1 score, on semantic score, on correctness score, and on completeness score, evaluated across three datasets. The most significant improvements were observed on LLM-as-judge scores over the LegalBench and SciDCC datasets. We also observed that all RAG methods perform much better than direct inference of LLMs without retrieval, as expected, showing the effectiveness of incorporating external knowledge for scientific QA tasks.
5.1.2. Retrieval performance.
Table 4.1 reports retrieval results for datasets with supporting passage annotations and models that explicitly retrieve passages. The retrieval accuracy of baseline methods also varied by dataset: HippoRAG 2 was the second-best model on SciFact, while e5 achieved this rank on both LegalBench and SciDCC. Compared to the strongest baselines, Hypercube-RAG exhibits superior precision and recall scores with an average improvement by and on SciFact, and on LegalBench, and and on SciDCC, respectively. These results highlight Hypercube-RAG as a state-of-the-art RAG system that enhances both retrieval and QA performance.
5.2. Efficiency
We study the response speed of diverse RAG methods over the corpus size and report the retrieval time for one query in Table 4.
5.2.1. Within corpus size .
@@ 5.2. Efficiency
BM25 achieves the fastest retrieval among all evaluated methods without surprise. In contrast, semantic and graph-based RAGs incur significantly higher retrieval costs. This is particularly pronounced for graph-based methods due to the computational burden of search paths in large-scale graph structures, posing scalability challenges. Notably, our Hypercube-RAG substantially reduces the retrieval time by one to two orders of magnitude compared to both semantic and graph-based methods, underscoring the effectiveness of the hypercube structure in optimizing retrieval efficiency. We attribute this to the retrieval with compact, multi-dimensional cube labels introduced in Section 3.3. Although it is slightly slower than BM25, the marginal increase in retrieval time is a worthwhile trade-off for significantly improved retrieval accuracy and response quality, as shown in Tables 4.1 and 4.1.
@@
5.2.2. Beyond corpus size .
To mimic the real-world scenario that the external knowledge database is usually large and noisy, we expand the hurricane-related corpus from 844 to 11,539 documents (around 14 times larger). The added documents are related to other topics, such as “Pollution” or “Ozone Holes”, introducing additional irrelevant documents (“noise”). Hypercube-RAG maintains the lowest retrieval time per query (see the last column in Table 4), highlighting the efficiency of the cube label-based retrieval mechanism. This is likely because most noisy documents result in empty or sparsely populated cube cells, as they lack corresponding fine-grained labels pre-defined in our hypercube. As a result, despite a significantly expanded corpus, the hypercube can efficiently bypass these irrelevant documents. In contrast, other baseline methods exhibit greater sensitivity to corpus size, resulting in increased retrieval times. In addition, Table 5 shows that the accuracy of the evaluated methods drops due to the inclusion of these noisy off-topic documents. However, Hypercube-RAG still outperforms others, demonstrating robust resilience to noise.
| Methods | Semantic Score with | Semantic Score with | Score () |
@@
5.3. Explainability
The retrieval process with our hypercube is inherently explainable through its associated labels in the cube cell(s). Considering the query “How much rainfall did Melbourne Beach, Florida receive from Tropical Storm Fay?” in Figure 3, Hypercube-RAG returns relevant documents by matching the decomposed components with fine-grained labels, “Melbourne Beach”, “Tropical Storm Fay”, and “rainfall”, along hypercube dimensions: Location, Event, and Theme. Table 6 shows that three retrieved documents (Doc. 565, 246, 535), which are represented with fine-grained cube labels. These cube labels clearly explain why those documents are returned. For example, Doc. 565 is ranked the first to return, since it contains all three key components of the query. The detailed document contents are included in Table 7.
@@ 5.3. Explainability
| Document ID | Document Content |
| 565 | …The 2008 Atlantic hurricane season started early with the formation of Tropical Storm Arthur on May 30, from the remnants of the eastern Pacific Ocean’s first storm, Alma, which crossed Central America and reformed in the Gulf of Mexico. It took one month and four days for the next storm to form, Bertha. Once a powerful Category 3 hurricane, now a tropical depression, Gustav moved from northwest Louisiana into northeastern Texas and into Arkansas by Sept. 3. Like Tropical Storm Fay in August, Gustav’s legacy will lie in large rainfall totals. According to the National Hurricane Center discussion on Sept. 2, ”Storm total rainfalls are expected to be five to ten Inches with isolated maximums of 15 inches over portions of Louisiana, Arkansas and Mississippi. Rainfall amounts of 4-8 inches have been already reported in parts of Alabama, Mississippi and Louisiana. …By the late morning, that tropical cyclone became Tropical Storm Josephine. In August, Fay’s ten-day romp from the U.S. Southeast northward up the Appalachian Mountains seemed like a harbinger for September’s storms. Fay took her time going northward and dumped tremendous amounts of rain along the way. Melbourne Beach, Florida, received as much as 25.28 inches of rain. Other cities in various states reported high totals: Thomasville, Ga., reported 17.43 inches; Camden, Ala., received 6.85 inches; Beaufort, S.C., received 6.11 inches; Carthage, Tenn., reported 5.30 inches, and Charlotte, N.C., reported 5.90 inches. Fay was a perfect example of how weaker tropical storms can cause flooding inland … |
| 535 | …Hopefully this research will aid the U.S. Army Corps of Engineers in identifying levees that need repair and assessing the lifespan of structures like the MRGO that are not planned for upkeep. Using fieldwork conducted in the Florida Panhandle, Straub and his fellow researchers were able to better understand the process of seepage erosion, which occurs when the re-emergence of groundwater at the surface shapes the Earth’s topography. The paper suggests that the velocity at which channel heads advance is proportional to the flux of groundwater to the heads. The researchers used field observations and numerical modeling to come up with the theory. … |
Retrieved Docs: [19, 230]
Doc 19: …Based on the ACE projection, combined with the above-average numbers of named storms and hurricanes, the likelihood of an above-normal Atlantic hurricane season has increased to 85%, with only a 10% chance of a near-normal season and a 5% chance of a below-normal season …, said Gerry Bell, Ph.D., lead seasonal hurricane forecaster at NOAA’s Climate Prediction Center …
Retrieved Docs: [230, 273]
Doc 230: …NOAA is predicting a 70 percent likelihood of 6 to 11 named storms, of which 3 to 6 could become hurricanes …While a below-normal season is likely (70 percent), there is also a 20 percent chance of a near-normal season, and a 10 percent chance of an above-normal season …
Retrieved Docs: [230, 636]
Doc 230: …NOAA is predicting a 70 percent likelihood of 6 to 11 named storms, of which 3 to 6 could become hurricanes …While a below-normal season is likely (70 percent), there is also a 20 percent chance of a near-normal season, and a 10 percent chance of an above-normal season …
Retrieved Docs: [583, 275]
Doc 583: …The Climate Prediction Center outlook calls for considerable activity with a 65 percent probability of an above normal season and a 25 percent probability of a near normal season. This means there is a 90 percent chance of a near or above normal season. …
5.4. Ablation Study
@@
Model Components. As shown in Section 3.3, Hypercube-RAG combines two retrieval strategies: sparse exact matching and dense embedding retrieval. Section 3.4 introduces the document ranking algorithm used to filter and prioritize the retrieved results. To study their effectiveness, we select the SciDCC dataset and conduct an ablation study by removing each of them from Hypercube-RAG, which are represented as No-Sparse, No-Dense, and No-Ranking, respectively. Figure 6 shows that the full version of Hypercube-RAG consistently outperforms all other variants, verifying positive contributions of each constituent. More specifically, it justifies the benefit of coupling these two strategies for scientific question-answering tasks, and the necessity of the ranking algorithm to filter possible “noisy” documents.
@@
Hypercube Dimensions. We also conduct an ablation study on the hypercube dimensions: Location, Event, Date, Organization, Person, and Theme. To study their effectiveness, we remove each of the dimensions, which are represented as No-Location, No-Event, No-Date, No-Organization, No-Person, and No-Theme, respectively. Table 8 shows that the full version of Hypercube-RAG consistently outperforms all other variants, verifying positive contributions of each dimension. We observe that the performance drops the most when removing the Location and Theme dimensions, justifying their high importance for theme-specific QA tasks in geoscience. We claim that these dimensions in a hypercube could be changed over datasets with different focuses of interest. The hypercube dimensions used for other datasets are provided in Appendix A.